OpenStudy (anonymous):

Write an equation for the line perpendicular to y = 5x – 12 that contains (–3, –1).

OpenStudy (solomonzelman):

even I can do that....

OpenStudy (luigi0210):

Use the point-slope formula

OpenStudy (luigi0210):

\(y-y_{1}=m(x-x_{1})\)

OpenStudy (luigi0210):

And since it's perpendicular, the slope \(m=5\) will be \(\frac{-1}{5}\)